OpenStudy (anonymous):

if |x| less than 0.5 then what is (4-8x)^-1/2??

OpenStudy (anonymous):

\[ |x|<0.5\implies -0.5<x<0.5 \]

OpenStudy (anonymous):

So our answer will also likely be in terms of a range.

OpenStudy (anonymous):

If you mutliply both sides by \(-8\) you get:\[ -4<-8x<4 \]

OpenStudy (anonymous):

Add \(4\) and you get: \[ 0<4-8x<8 \]

OpenStudy (anonymous):

Take square root and you get: \[ 0<\sqrt{4-8x}<\sqrt 8 \]

OpenStudy (anonymous):

Finally, you invert it and you get :\[ \frac{1}{\sqrt 8} < \frac{1}{\sqrt{4-8x}} \]

OpenStudy (anonymous):

it is not an answer(

OpenStudy (anonymous):

Sure it is!

OpenStudy (anonymous):

x|<0.5⟹−0.5<x<0.5 it is not true)))

OpenStudy (anonymous):

\[ |x| < 0.5 \]means that: \[ x < 0.5 \] and \[ -x < 0.5 \]We can simplify \(-x < 0.5\): multiply by -1: \(x > -0.5\) which is just \(-0.5 < x\) If you combine \(-0.5 <x \) and \(x<0.5\) you get \(-0.5<x<0.5\).

OpenStudy (anonymous):

ok